A fly wheel slows from 600 to 400rev/min while rotating through 40 revolutions.
A, what is the angular acceleration of the fly wheel?
B, how much time elapses during the 40 revolutions?
1
Expert's answer
2020-05-11T20:11:00-0400

a)


"\omega^2-\omega^2_0=2\alpha \theta\to 400^2-600^2=2(40)\alpha"

"\alpha=-2500\frac{rev}{min^2}"

b)


"400-600=-2500t"

"t=0.08\ min"